### v1

# Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning

Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that maximizing the mutual information between the task and the task representation ($I(Z;M)$) can lead to performance improvements. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. Motivated by the return discrepancy scheme in the model-based RL field, we find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, this optimization process ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound to consider the impacts of task representation shift explicitly. We demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against previous approaches. To make it practical, we design an easy yet highly effective algorithm RETRO (REining Task Representation shift in context-based Offline meta reinforcement learning) with only adding one line of code compared to the backbone. Empirical results validate its state-of-the-art (SOTA) asymptotic performance, training stability and training-time consumption on MuJoCo and MetaWorld benchmarks.
1 Introduction
Reinforcement Learning (RL) has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering], primarily through online RL methods. However, the extensive interactions of online RL entail high costs and safety concerns, posing significant challenges in real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting, li2024towards] has emerged as a beacon to illuminate the way toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Along the line of OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] wield a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, they propose to train a context encoder independently to learn the task representation from the collected context (e.g. trajectories), then pass the latent representation to the downstream policy and value function, forming a decoupled framework. Thanks to [li2024towards] for unveiling the nature of the prior COMRL works that they predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements [li2020focal, yuan2022robust, gao2024context, li2024towards]. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
Intriguingly, we unify the mutual information maximization objective commonly used in COMRL with the general RL objective of maximizing expected return. To be specific, maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee, filling the gap between intuition and theoretical justification. However, after scrutinizing this optimization process, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we adopt the framework of performance difference bound to consider the impacts of task representation shift explicitly. We theoretically demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against prior works.
To make it practical, we derive an optimization problem concerning task representation shift and find that reining the update frequency of task representation and employing a straightforward classifier-based method for task representation extraction is sufficient to beat existing SOTA algorithms. RETRO is evaluated on a wide range of RL benchmarks covering MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta]. With the addition of just one line of code, RETRO achieves superior asymptotic performance and training stability compared to the classifier-based method alone. Notably, RETRO also holds a tuning-free property. Moreover, our approach to reining task representation shift is general and is also ablated in the experiments.
2 Related Works
Offline RL. Offline RL [fujimoto2019off] leverages a fully offline dataset to perform policy optimization and value function learning, avoiding explicit interaction with the environment. To overcome the notorious distribution shift and the induced value overestimation issues, prior works adopt methods based on policy regularization [fujimoto2019off, kumar2019stabilizing, fujimoto2021minimalist], value penalty [wu2019behavior, kumar2020conservative] and sequence modeling [chen2021decision, janner2021offline]. While these methods have demonstrated promising performance on a single task, their generalization performance to the unknown remains limited [ghosh2021generalization]. Our work follows the offline RL setting, attempting to adopt meta-learning techniques to improve the generalization performance.
Meta RL. Meta RL has emerged as an enticing avenue for effectively adapting to unknown environments [finn2017model]. Unlike traditional gradient-based meta RL methods [finn2017model, mitchell2021offline], context-based meta RL methods demonstrate superior advantages in both sample efficiency and asymptotic performance [rakelly2019efficient, zintgraf2019varibad, li2024towards]. Our work delves into this promising domain to advance the capabilities of context-based meta RL further.
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of learned task representation to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], many prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements. Though achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. We introduce the return discrepancy scheme [janner2019trust] into the COMRL setting to provide a feasible explanation. However, after scrutinizing this optimization process, we further find that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. Centered around this motivation, our work aims to present a theoretical framework that reins the task representation shift.
Performance Improvement Guarantee. Performance improvement guarantee is a fundamental concern in both online RL and offline RL settings. In online RL, performance difference bound [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic] and regret bound [osband2014model, curi2020efficient] are widely adopted schemes to get a performance improvement guarantee. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the multi-task and COMRL settings largely unexplored. We first provide a theoretical justification w.r.t the performance improvement guarantee for prior COMRL works. Then, we prove that it is possible to achieve monotonic performance improvements by reining the task representation shift ignored by prior COMRL works, thereby showcasing the advantage against them.
3 Preliminaries
3.1 Problem Statement
A task in RL is formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$. A viable solution is to iterate the Bellman operator over the action-value function [sutton2018reinforcement], which is denoted as $Q_{\pi}(s,a)=\mathbb{E}_{\pi,P}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
It is widely assumed that the task $m$ in the meta RL setting is subject to the distribution $p(M)$ with sharing the same state space and action space but different from transition dynamics and reward functions [li2020focal, yuan2022robust, gao2024context, li2024towards]. At training time, tasks are randomly sampled from the task distribution $p(M)$. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with the only right to access the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
3.2 Context-based OMRL
COMRL holds a steadfast belief that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, COMRL methods would train a task encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exist the optimal task representation distribution $Z^{*}(\cdot|x)$ that can deduce the optimal expected return. The objective of COMRL is formalized as:
$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
where $R_{m}(s,a)$ denotes the ground-truth reward function of the task $m$.
Prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvement, as shown in Theorem 3.1.
Theorem 3.1 ([li2024towards]).
Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. When meeting $I(Z;M;X_{b})\geq 0$:
$I(Z;X_{t}|X_{b})\leq I(Z;M)\leq I(Z;X)$ (2)
where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
However, the theoretical justification of performance improvement for such intuition has been lacking.
4 Methods
To fill the gap, we introduce the return discrepancy scheme [janner2019trust] into the COMRL setting and find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation (Section 4.1). Further, after scrutinizing this optimization process, we argue that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. We then turn to use the framework of performance difference bound to consider task representation shift explicitly. We theoretically prove that it is possible to achieve monotonic performance improvements by reining task representation shift, thereby showcasing the advantage against prior works (Section 4.2). To make it practical, we derive an optimization problem concerning task representation shift and use the simplest algorithm to instantiate it (Section 4.3).
We first introduce the Lipschitz assumption for the following use. Lipschitz-based assumptions are widely adopted in both model-free RL [song2019efficient, ghosh2022provably] and model-based RL [ji2022update, zhang2023how].
Assumption 4.1.
The policy function is $L_{z}-$ Lipschitz with respect to some norm |||||$\cdot$||| | ⋅ | | in the sense that
$\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to induce an optimization objective served for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
Definition 4.2 (Return discrepancy in COMRL).
The return discrepancy in the COMRL setting can be defined as
$J^{*}(\theta)-J(\theta)\geq-|J(\theta)-J^{*}(\theta)|$ (4)
where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the task encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage iteration framework with the first stage optimizing the task encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage optimizing the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, which forms a bedrock to connect the bridge between the intuition of prior works and the theoretical justification concerning the performance improvement guarantee.
Theorem 4.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
Proof: See Section 7.2 in the Appendix. ∎
By directly unrolling the Theorem 4.3, we can establish the lower bound of $J^{*}(\theta)$ as:
$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (6)
Note that $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$ is a constant w.r.t the learned task representation $Z(\cdot|x)$. Hence, this term can be ignored when optimizing $Z(\cdot|x)$. Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$, but this is not our main focus and we leave this to future work.
Theorem 4.3 and Eq. (6) indicate the following two principles for the task encoder and the policy. For the task encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation is irrelevant for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that prior COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$ and training the policy network by conditioning on the learned task representation as well as applying the standard offline RL algorithms. As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the task representation shift between two consecutive updates alone. In the next section, we will show that the impacts of task representation shift may lead to performance improvement collapse.
4.2 Monotonic Performance Improvement Concerning Task Representation Shift
We attempt to build a theoretical framework that can achieve monotonic performance improvement. Hence, we turn to use the framework of performance difference bound.
Definition 4.4 (Performance difference bound in COMRL).
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the task encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the task encoder and the policy. The performance difference bound in the COMRL setting can be defined as
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq C$ (7)
when $C$ is non-negative, the algorithm allows a monotonic performance improvement in the real environment.
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we transform Eq. (7) to the following formulation, where $J^{mutual}(\theta_{i}),i\in[1,2]$ denote the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks.
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ $=(J^{2}(\theta_{2})-J^{mutual}(\theta_{2}))+(J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1}))$
$+(J^{mutual}(\theta_{1})-J^{2}(\theta_{1}))+(J^{2}(\theta_{1})-J^{1}(\theta_{1}))$ (8)
Then, we introduce a mild assumption for the task representation distribution.
Assumption 4.5.
The support set of $Z_{i}(\cdot|x),i\in[1,2]$ covers the support set of $Z_{mutual}(\cdot|x)$.
Assumption 4.5 usually holds in practice as when we give $z\sim Z_{mutual}(\cdot|x)$ to the task representation distribution $Z_{i}(\cdot|x)$ parameterized by the neural network, we can always get a non-zero probability. Given that augmented state $s^{+}=[s,z]$, by applying standard offline RL algorithms with enough samples, the policy update would lead to performance improvement when meeting Assumption 4.5, which means
$\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (9)
By leveraging Eq. (4.2) and the above assumptions, we derive the lower bound of the performance difference.
Theorem 4.6 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (10)
Proof: See Section 7.2 in the Appendix. ∎
For the need of monotonic performance improvement guarantee, we further let the right-hand side of Eq. (10) be positive. Then, we would have the following requirement to achieve monotonicity.
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (11)
The above equation unveils the key influence of the task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If the task representation shift is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (11). On the contrary, if the task representation shift is too large, the right-hand side of Eq.(11) may be negative. Both of these cases would lead to performance improvement collapse. Thus, we argue that the impacts of task representation shift cannot be ignored.
To give an example for this problem, we introduce the following assumptions.
Assumption 4.7.
The impacts of task representation shift can be upper-bounded by $\beta$.
Assumption 4.8.
The space of the task representation is discrete and limited.
Assumption 4.9.
The i.i.d samples from $Z_{mutual}(\cdot|x)$ can be obtained.
Assumption 4.10.
The task encoder is trained by fitting an empirical distribution on i.i.d samples from $Z_{mutual}(\cdot|x)$.
Though sampling from $Z_{mutual}(\cdot|x)$ seems impractical, we can approximate it as sampling from a task encoder pre-trained by various bounds of $I(Z;M)$.
Corollary 4.11 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (12)
Here, $\xi\in[0,1]$ is a constant.
Proof: See Section 7.2 in the Appendix. ∎
4.3 Practical Implementation
To simplest make it practical, we first transform Eq. (11) to an optimization problem.
$\min|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|,\,\,s.t.\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|>0$ (13)
As alluded in Section 4.1, many contrastive learning-based algorithms involved in prior works [li2020focal, yuan2022robust, gao2024context, li2024towards] can be adopted to minimize $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$. Unfortunately, no one among them is free from parameter-tuning. We propose to use a classifier-based manner to address this concern. Specifically, after extracting from the task encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the task encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. Note that this classifier-based manner enjoys tuning-free property and also can be seen as an approximation of $I(Z;M)$. To the best of our knowledge, we are the first to propose adopting a classifier-based manner to train the task encoder in the COMRL setting.
For reining task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, since the values lying in the constraint of Eq. (13) are intractable, approximately estimating them may introduce a series of hyper-parameters thus significantly increasing the complexity of the algorithm. Instead, one insight from Corollary 4.11 is that tuning the update of the task encoder may bring better performance improvements. Through the experiment, we find that adjusting the update frequency of the task encoder to half of the origin is sufficient to beat the existing SOTA algorithms. Empirical results demonstrate that the previously ignored task representation shift does make a remarkable difference.
Our main algorithm is shown in Algorithm 1. Compared to the backbone, RETRO only adds one line of code, which is colored by red. To maintain the consistency of prior COMRL works, we also adopt BRAC [wu2019behavior] to train the policy. Moreover, combining prior works with our approach to reining task representation shift is also ablated in the experiments.
Algorithm 1 RETRO Framework
0:  Offline Datasets $\bm{X}=\{x^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{m^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\theta}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, task batch size $N_{tbatch}$
1:  while not done do
2:     Sample task $\{m^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{x^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
3:     for step in each iteration do
4:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
5:        Obtain task representations $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
6:        // Rein the task representation shift
7:        if step % 2 == 0 then
8:           Compute $\mathcal{L}_{\text{RETRO}}=\textit{CrossEntropyLoss}(z^{j},m^{j})$
9:           Update $\phi$ to minimize $\mathcal{L}_{\text{RETRO}}$
10:        end if
11:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
12:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
13:        Compute $\mathcal{L}_{\text{BRAC}}$ to update $\theta,\psi$
14:     end for
15:  end while
5 Experiment
Our experiments are conducted to answer 4 questions: 1) Compared to existing SOTA algorithms, can RETRO achieve more stable performance improvements and higher asymptotic performance? 2) Can RETRO be integrated into prior works like UNICORN and FOCAL? 3) How does reining the task representation shift with different frequencies affect RETRO? 4) Compared to existing SOTA algorithms, can RETRO achieve lower training-time consumption?
5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. We choose 4 competitive baselines to validate our superiority. The baselines cover UNICORN [li2024towards], which adopts reconstruction loss to balance the impacts of spurious correlation, CSRO [gao2024context], which adopts CLUB loss to upper-bound the impacts of the behavior policy, FOCAL [li2020focal], which directly applies distance metric learning to train the task encoder. and Classifier, which is our backbone without reining the task representation shift. The implementation details of all algorithms are shown in Appendix Section 7.3. Note that all these baselines can be categorized into the return discrepancy scheme stated in Section 4.1. Following the protocol of UNICORN, we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
5.2 Baseline Comparisons
Figure 1: Testing returns of RETRO against baselines on six benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 1 illustrates the learning curves of RETRO against other baselines on 20 testing tasks for each benchmark. The results showcase that RETRO can unanimously achieve SOTA asymptotic performance across all benchmarks. Compared to the backbone Classifier, with adding only one line of code to rein the task representation shift, RETRO achieves remarkable performance improvements (e.g. approximately 20% relative performance improvements on Ant-Dir and 25% relative performance improvements on Dial-Turn). It is worth noting that RETRO shows comparably stable performance improvements in all the benchmarks especially on Point-Robot and Dial-Turn. Recall that RETRO only needs to adjust the update frequency to half of its origin, thus leading to a tuning-free algorithm. On the contrary, both UNICORN and CSRO are especially sensitive to the hyper-parameter settings. This phenomenon is particularly evident in the extremely challenging benchmark Humanoid-Dir, which holds the dimension of state space 376. We argue that without carefully parameter-tuning trials, UNICORN and CSRO cannot exhibit the desired performance. Admittedly, though more advanced algorithms may exist to better control the task representation shift, adopting the simplest manner to achieve SOTA performance makes RETRO more applicable to other fields.
5.3 Can RETRO Be Integrated Into Other Methods?
According to our analysis in Section 4.1, prior COMRL works ignore the impacts of task representation shift, which may lead to performance collapse. As our main focus is to rein the task representation shift, which is orthogonal to the prior works, RETRO should bring better performance improvements when being integrated into prior works. We denote RETRO-UNICORN and RETRO-FOCAL as combining the reining method into the original UNICORN and FOCAL respectively. To ensure fairness, we also adjust the update frequency of the task encoder in RETRO-UNICORN and RETRO-FOCAL to half of the origin.
Figure 2: Testing returns of RETRO-UNICORN and RETRO-FOCAL against the origin on three benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 demonstrates the learning curves of RETRO-UNICORN and RETRO-FOCAL against their backbones on 20 testing tasks for each benchmark. The results illustrate that adopting RETRO enables better performance improvements across all the benchmarks, thereby validating our core claim on reining the impacts of task representation shift. Note that RETRO(-Classifier) has comparable or even higher performance than RETRO-UNICORN, we believe that RETRO(-Classifier) should be a preferred option due to the tuning-free property.
Though our goal is to design the simplest algorithm to prove the effectiveness of reining the task representation shift, we can also notice that the convergence speed is relatively lower than the backbones at the beginning of the training process. Hence, in the future, we envision designing a strong algorithm to control the task representation shift delicately, where the springboard is RETRO.
5.4 The Influence Of Different Update Frequencies
Figure 3: Testing returns of difference update frequency on two benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 4: Average training time consumption of RETRO against other baselines on six benchmarks.
Figure 4 illustrates the influence of different update frequencies on the performance of RETRO. The results show that setting the update frequency to half of the origin can generally strike a balance among the convergence speed, training stability and asymptotic performance across all the benchmarks. Hence, we argue that adopting RETRO as a tuning-free algorithm is acceptable.
5.5 Can RETRO Reduce Training-Time Consumption?
We report the training-time consumption of RETRO against other baselines across all benchmarks in Figure 4. All algorithms are trained with the same architecture on two RTX3090 graphic cards. With the exception of Humanoid-Dir, which can only run one seed per graphic card due to the limitation, the reported times for the remaining benchmarks pertain to running two seeds per card and a total of four seeds concurrently during training. Figure 4 shows that RETRO has a significant time-consumption reduction against UNICORN [li2024towards], CSRO [gao2024context] and FOCAL [li2020focal]. Though this fact is based on our choice of backbone to be a classifier, it is the one proposed for the first time in our paper aiming at achieving a tuning-free algorithm. Compared to the backbone classifier, RETRO also achieves lower training-time consumption, as RETRO reduces the update frequency of the task encoder.
6 Conclusion & Limitation
Finding that the theoretical justification of maximizing $I(Z;M)$ can lead to performance improvement has been lacking, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, after scrutinizing this optimization framework, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound, proving that by reining the task representation shift, it is possible to achieve monotonic performance improvements. To make it practical, we propose a tuning-free algorithm RETRO with only adding one line of code compared to the backbone. Extensive experiments have validated the superiority of RETRO on training stability, asymptotic performance and training-time consumption. Though we choose classifier-based loss to train the task encoder, the idea that reins the task representation shift is general, which is shown in Section 5.3. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift, which is also stated in Section 5.3.
[…]
7 Appendix
7.1 Useful Lemmas
Lemma 7.1 (Return bound.[zhang2023how]).
Let $R_{max}$ denote the bound of the reward function, $\epsilon_{\pi}$ denote $\max_{s}D_{TV}(\pi_{1}||\pi_{2})$ and $\epsilon^{M_{2}}_{M_{1}}$ denote $\mathbb{E}_{(s,a)\sim d^{\pi_{1}}_{M_{1}}}[D_{TV}(p_{M_{1}}||p_{M_{2}})]$. For two arbitrary policies $\pi_{1},\pi_{2}\in\Pi$, the expected return under two arbitrary models $M_{1},M_{2}\in\mathcal{M}$ can be bounded as,
$|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|\leq 2R_{max}(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (14)
Lemma 7.2 (Inequility for $L_{1}$ deviation of the empirical distribution.[weissman2003inequalities]).
Let $P$ be a probability distribution on the set $\mathcal{A}=\{1,...,a\}$. For a sequence of samples $x_{1},...,x_{m}\sim P$, let $\hat{P}$ be the empirical probability distribution on $\mathcal{A}$ defined by $\hat{P}(j)=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(x_{i}=j)$. The $L_{1}$-deviation of the true distribution $P$ and the empirical distribution $\hat{P}$ over $\hat{A}$ from $m$ independent identically samples is bounded by,
$Pr(|P-\hat{P}|\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp{(-m\epsilon^{2}/2)}$ (15)
7.2 Missing Proofs
Theorem 7.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (16)
Proof:
$|J(\theta)-J^{*}(\theta)|$
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]-\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (17)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (18)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]$
$+\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (19)
$\leq\mathbb{E}_{m,x}[|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]|$
$+|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]|]$ (20)
$=\mathbb{E}_{m,x}[|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|$
$+|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|]$ (21)
$\leq\mathbb{E}_{m,x}[\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta)||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta))$
$+\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)||\pi(\cdot|s,Z^{*}(\cdot|x);\theta))]$ (22)
$\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (23)
Eq. (22) is the result of directly applying Lemma 7.2 and Eq. (23) is the result of directly applying Assumption 4.1. ∎
Theorem 7.4 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (24)
Proof:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ (25)
$=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (26)
$=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]]$ (27)
Denote $\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{i}(\cdot|x);\theta_{j})}}[R_{m}(s,a)]]$ as $G^{\theta_{j}}_{Z_{i}}$, then we can get the following derivation. Here, we only consider the value within the brackets.
$=\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (28)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (29)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (30)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 7.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (31)
$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))+\epsilon^{mutual}_{12}$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1}))$ (32)
$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}$
$-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (33)
Simplifying Eq. (33) we can get:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (34)
Corollary 7.5 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (35)
Here, $\xi\in[0,1]$ is a constant.
Proof: Let $Z_{1}$ has already been trained on $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ to get the monotonic performance improvement guarantee.
By applying Lemma 7.2, the $L_{1}$ deviation of the empirical distribution $Z_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{|\mathcal{\mathbb{Z}}|}-2)\exp{(-(N+k)\epsilon^{2}/2)}$ (36)
Then for a fixed $x$, with probability greater than $1-\xi$, we have:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{N+k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (37)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (38)
7.3 Implementation Details
In this section, we report the implementation details of RETRO and all the baselines.
  • FOCAL[li2020focal]. We adopt the open-source code of FOCAL [^1] https://github.com/FOCAL-ICLR/FOCAL-ICLR/. To better show the performance of FOCAL in each environment, we tune the parameter $\beta$ in distance metric learning from $[0.1,0.5,1]$ and choose the best result in our baseline comparison.
  • CSRO[gao2024context]. We adopt the open-source code of CSRO [^2] https://github.com/MoreanP/CSRO. To better show the performance of CSRO in each environment, we tune the club loss weight and task representation loss weight from $[10,20,30]\times[5,10,15]$ and choose the best result in our baseline comparison.
  • UNICORN[li2024towards]. We obtain the source code of UNICORN from the authors. To better show the performance of UNICORN in each environment, we tune the parameter $\frac{\alpha\beta-1}{(1-\alpha\beta)}$ from $[0.1,0.3,0.7,1,1.5]$ and choose the best result in our baseline comparison.
  • Classifier. Built upon the code base of FOCAL, Classifier only replaces the distance metric learning loss with cross-entropy loss, achieving tuning-free property.
  • RETRO. Built upon the code base of Classifier, RETRO only adjusts the update frequency of the task encoder to half of the origin, by adding merely one row of code.
To maintain fairness, we make sure all benchmark-irrelevant parameters are consistent, as shown in Table 1.
Table 1: Benchmark-irrelevant parameters setting in the training process.
training tasks 20
testing tasks 20
task training batch size 16
rl batch size 256
context size 1 trajectory
actor-network size [256, 256]
critic-network size [256, 256]
task encoder network size [64, 64]
learning rate 4e-3
Table 2: Benchmark-relevant parameters setting in the training process.
Configurations Point-Robot Humanoid-Dir Reach Ant-Dir Button-Press Dial-Turn
dataset size 2100 9e4 3e5 9e4 3e5 3e5
task representation dimension 5 64 5 5 5 5
For each environment, we make the benchmark-relevant parameters the same for each algorithm, as shown in Table 2.