### v2

# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task and the task representation ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored. Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby providing a feasible explanation concerning performance improvements. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating optimization process, which may lead to performance improvement collapse. We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates. We set different manners to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
1 Introduction
RL has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering, zhang2023safe], primarily through online RL methods. However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting] has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Among the OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, these methods [li2020focal, gao2024context, li2024towards] propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g. trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left. The context encoder and policy are trained in an alternating manner, with each being updated once per cycle. Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
Figure 1: Our training framework compared to the previous training framework. They both adopt the alternating optimization framework to train the context encoder and the policy. However, ours considers the variation of task representation ignored by the previous training framework.
Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return. To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process. Variations of the task representation that are either too large or too small can lead to performance improvement collapse. To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement guarantees with proper updates.
To show the impacts of this issue, we set different manners to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities. These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards]. We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$. Empirical results show that reining in the task representation shift can indeed improve performance. Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
2 Related Works
Offline RL. Offline RL [fujimoto2019off] leverages a fully offline dataset to perform policy optimization and value function learning, avoiding explicit interaction with the environment. To overcome the notorious distribution shift and the induced value overestimation issues, prior works adopt methods based on policy regularization [fujimoto2019off, kumar2019stabilizing, fujimoto2021minimalist], value penalty [wu2019behavior, kumar2020conservative] and sequence modeling [chen2021decision, janner2021offline]. While these methods have demonstrated promising performance on a single task, their generalization performance to the unknown remains limited [ghosh2021generalization]. Our work follows the offline RL setting, attempting to adopt meta-learning techniques to improve the generalization performance.
Meta RL. Meta RL has emerged as an enticing avenue for effectively adapting to unknown environments [finn2017model]. Unlike traditional gradient-based meta RL methods [finn2017model, mitchell2021offline], context-based meta RL methods demonstrate superior advantages in both sample efficiency and asymptotic performance [rakelly2019efficient, zintgraf2019varibad, li2024towards]. Our work delves into this promising domain to advance the capabilities of context-based meta RL further.
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$. However, our work moves the focus on how to get better asymptotic performance based on a given context encoder learning algorithm. Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
Performance Improvement Guarantee.  Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings. In online RL, performance improvement guarantees are often established through methods such as performance difference bounds [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic], and regret bounds [osband2014model, curi2020efficient]. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the COMRL settings largely unexplored. While ContrBAR [choshen2023contrabar] also benefits from the performance improvement guarantee in the OMRL setting, it is specifically served for its own framework. In contrast, our work addresses a broader class of algorithms that maximize $I(Z;M)$, including but not limited to ContrBAR.
3 Preliminaries
3.1 Problem Statement
A task in RL is generally formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
Given a distribution of tasks $p(M)$, every task $m$ sampled from $p(M)$ is an MDP sharing the same state space and action space but differing from transition dynamics and reward functions. It is widely assumed that the tasks in the meta RL setting are randomly sampled from $p(M)$ [li2020focal, yuan2022robust, gao2024context, li2024towards]. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with access only to the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
3.2 Context-based OMRL
Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$. Hence, the objective of COMRL is formalized as:
$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
where $R_{m}(s,a)$ denotes the ground-truth reward function of the task $m$.
Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1. They hold the intuition that using such kind of manner to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements. However, the theoretical justification of performance improvement for such intuition has been less explored.
Theorem 3.1 ([li2024towards]).
Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. We have:
$I(Z;X_{t}|X_{b})\leq I(Z;M)\leq I(Z;X)$ (2)
where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
4 Methods
We first introduce a Lipschitz assumption, a widely used technique in both model-free [song2019efficient, ghosh2022provably] and model-based [ji2022update, zhang2023how] RL frameworks.
Assumption 4.1.
The policy function is $L_{z}-$ Lipschitz w.r.t some norm |||||$\cdot$||| | ⋅ | | in the sense that
$\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to induce an optimization objective served for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
Definition 4.2 (Return discrepancy in COMRL).
The return discrepancy in the COMRL setting can be defined as
$J^{*}(\theta)-J(\theta)\geq-|J(\theta)-J^{*}(\theta)|$ (4)
where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the context encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, connecting the bridge between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
Theorem 4.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ as the optimal task representation distribution served for maximizing $I(Z;M)$  [^1] Whether the solution of maximizing $I(Z;M)$ is equivalent to the solution of maximizing the expected return is still an open problem. Hence, we introduce an extra notation here.. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
Proof: See Section 8.2 in the Appendix. ∎
By directly unrolling the Theorem 4.3, we can establish the lower bound of $J^{*}(\theta)$ as:
$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (6)
Note that $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$ is a constant w.r.t the learned task representation $Z(\cdot|x)$. Hence, this term can be ignored when optimizing $Z(\cdot|x)$. Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$, but this is not our main focus and we leave this to future work.
Theorem 4.3 and Eq. (6) indicate the following two principles for the context encoder and the policy. For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation related terms are constants for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that prior COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$ and training the policy network by conditioning on the learned task representation as well as applying the standard offline RL algorithms. As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the variation of the task representation alone. In the next section, we will show that this may lead to performance improvement collapse.
4.2 Monotonic Performance Improvement Concerning Task Representation Shift
We attempt to build a theoretical framework that can achieve monotonic performance improvement. Hence, we turn to use the framework of performance difference bound.
Definition 4.4 (Performance difference bound in COMRL).
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the context encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the context encoder and the policy. The performance difference bound in the COMRL setting can be defined as
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq C$ (7)
when $C$ is non-negative, the algorithm allows a monotonic performance improvement.
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. We denote $J^{mutual}(\theta_{i}),i\in[1,2]$ as the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks and then introduce a mild assumption for the task representation distribution.
Assumption 4.5.
The support set of $Z_{i}(\cdot|x),i\in[1,2]$ covers the support set of $Z_{mutual}(\cdot|x)$.
Assumption 4.5 usually holds in practice as when we give $z\sim Z_{mutual}(\cdot|x)$ to the task representation distribution $Z_{i}(\cdot|x)$ parameterized by the neural network, we can always get a non-zero probability. Given that augmented state $s^{+}=[s,z]$, by applying standard offline RL algorithms with enough samples, the policy update would lead to performance improvement when meeting Assumption 4.5, which means
$\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (8)
By leveraging Definition 4.4 and the above assumptions, we derive the lower bound of the performance difference.
Theorem 4.6 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of the performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
Proof: See Section 8.2 in the Appendix. ∎
For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (9) be positive. Then, we would have the following condition to achieve monotonicity.
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (10)
The above equation unveils the key influence of the variation of task representation $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (10). On the contrary, if $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too large, the right-hand side of Eq.(10) may be negative. Both of these cases would lead to the performance improvement collapse. Thus, we argue that the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ cannot be ignored.
To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue task representation shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement on the condition of proper context encoder updates, as shown in Theorem 4.10.
Assumption 4.7.
The impacts of task representation shift can be upper-bounded by $\beta$.
Assumption 4.8.
The space of the task representation is discrete and limited.
Assumption 4.9.
There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ from the training dataset based on $Z_{1}(\cdot|x)$ and $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^2] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $Z^{empirical}_{2}$ should become similar..
Theorem 4.10 (Monotonic performance improvement guarantee on training process).
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
Here, $\xi\in[0,1]$ is a constant.
Proof: See Section 8.2 in the Appendix. ∎
Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{mutual}_{12}$. If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{mutual}_{12}$. As $\epsilon^{mutual}_{12}$ increases, $k$ decreases. When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder. Notice that this updating process can be performed multiple times, as long as satisfying Assumption 4.7. The updating process is visualized in Figure 1 right. Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works. The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.
Algorithm 1 General Algorithmic Framework Towards Reining In The Task Representation Shift
0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $q_{\phi}$ and given task representation shift threshold $\beta$
1:  for iter in alternating iterations do
2:     // Update the context encoder
3:     if Conditions for updating the context encoder are met then
4:        while Accumulated task representation shift is less than $\beta$ do
5:           Sample context from $\bm{X}$
6:           Obtain task representations from $q_{\phi}$ with inputting the sampled context
7:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
8:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
9:        end while
10:     end if
11:     // Update the policy
12:     Detach the task representations
13:     Sample training data from $\bm{X}$
14:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
15:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
16:  end for
4.3 Practical Implementation
To show the impacts of the task representation shift issue, we set different manners to adjust the context encoder. We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$. Our manners cover 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$. All manners as well as the original manner, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors. The reconstruction-based algorithm, passing the state, action, and task representation through the decoder to reconstruct the next state and the reward signal, is applied in [zintgraf2019varibad, dorfman2021offline, li2024towards] and is proven to be the lower bound of $I(Z;M)$. The contrastive-based algorithm, using the task representation to compute distance metric learning loss, is applied in [li2020focal, gao2024context, li2024towards] and is proven to be the upper bound of $I(Z;M)$ as well as contain the lower bound of $I(Z;M)$. The cross-entropy-based algorithm is a straightforward discrete approximation to $I(Z;M)$ and is proposed by our work. Specifically, after extracting from the context encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the context encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. The implementation details of these algorithms are shown in Appendix Section 8.3.
With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC [wu2019behavior] to train the policy.
5 Experiment
Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient. We hope to illustrate the necessity of reining in the task representation shift, laying the foundation for further research towards better asymptotic performance of COMRL.
5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. Following the protocol of UNICORN  [li2024towards], we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
5.2 MAIN RESULTS
Figure 2: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 illustrates the learning curves of our different manners to rein in the task representation shift against the original manner on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$. We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance. These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$. While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential. As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
Figure 3: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$ on different data qualities of Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
5.3 Can The Results Show Consistency Across Different Data Qualities?
To make our claim more universal, we conduct the experiment on different data qualities. We collect three types of datasets with size being equal to the dataset used in Section 5.2. Depending on the quality of the behavior policies, we denote these datasets as random, medium, and expert respectively.
Figure 3 demonstrates the learning curves of our different manners to rein in the task representation shift against the original manner concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
Based on these impressive results, we believe that developing a smarter algorithm to control the task representation shift automatically is an appealing direction and we leave this to future work.
6 Discussion
This section seeks to raise some interesting issues derived from our work. We hope these issues can further enable pondering on the significance of the task representation shift.
Figure 4: Testing returns of the pre-training scheme against training from scratch of cross-entropy, contrastive, and reconstruction based objectives on Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
According to our analysis, the task representation shift issue only happens in the case that the context encoder needs to be trained from scratch. Hence, a question is raised naturally. If we train the context encoder in advance and use this pre-trained context encoder to train the policy directly, can we achieve better performance improvement?
To answer this question, we conduct the experiment on Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder. As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch. To better theoretically interpret this, we introduce the following Corollary.
Corollary 6.1 (Monotonic performance improvement condition for pre-training scheme).
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (12)
Proof: See Section 8.2 in the Appendix. ∎
According to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements as $|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|$ is not equal to 0. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $I(Z;M)$. Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts stemming from the variation of the task representation in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
Figure 5: The 2D projection of the learned task representation space in Ant-Dir. Points are uniformly sampled from the evaluation datasets. Tasks of given goals from 0 to 6 are mapped to rainbow colors, ranging from purple to red.
6.2 Can The Visualization Of Task Representation Be Strongly Relied Upon to Imply The Asymptotic Performance?
Previous COMRL endeavors [li2020focal, yuan2022robust, gao2024context] mostly apply t-SNE on the learned task representation to show the differentiation of the tasks. They hold the belief that better performance corresponds to better differentiation of the tasks. However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation variation during the whole optimization process. According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance. Hence, it is insufficient to imply performance based on such evaluation principles. Nevertheless, we recognize that the visualization result can be seen as an auxiliary metric to assist in determining the task representation.
7 Conclusion & Limitation
Finding that the performance improvement guarantee of the COMRL training framework remains underexplored, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that the previous optimization framework can be linked with the general RL objective of maximizing the expected return. Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation. Based on this finding, we propose a new issue called task representation shift and prove that it is possible to achieve monotonic performance improvements with proper context encoder updates. With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift.
[…]
8 Appendix
8.1 Useful Lemmas
Lemma 8.1 (Return bound.[zhang2023how]).
Let $R_{max}$ denote the bound of the reward function, $\epsilon_{\pi}$ denote $\max_{s}D_{TV}(\pi_{1}||\pi_{2})$ and $\epsilon^{M_{2}}_{M_{1}}$ denote $\mathbb{E}_{(s,a)\sim d^{\pi_{1}}_{M_{1}}}[D_{TV}(p_{M_{1}}||p_{M_{2}})]$. For two arbitrary policies $\pi_{1},\pi_{2}\in\Pi$, the expected return under two arbitrary models $M_{1},M_{2}\in\mathcal{M}$ can be bounded as,
$|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|\leq 2R_{max}(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (13)
Proof:
$\begin{split}|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|&=|\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}(p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a))r(s,a)|\\ &\leq R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}|p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a)|\\ &=2R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\\ \end{split}$ (14)
According to Theorem 2 (Return Bound) in [zhang2023how],
$\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\leq(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (15)
We bring this result back and can draw the conclusion.
Lemma 8.2 (Inequility for $L_{1}$ deviation of the empirical distribution.[weissman2003inequalities]).
Let $P$ be a probability distribution on the set $\mathcal{A}=\{1,...,a\}$. For a sequence of samples $x_{1},...,x_{m}\sim P$, let $\hat{P}$ be the empirical probability distribution on $\mathcal{A}$ defined by $\hat{P}(j)=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(x_{i}=j)$. The $L_{1}$-deviation of the true distribution $P$ and the empirical distribution $\hat{P}$ over $\hat{A}$ from $m$ independent identically samples is bounded by,
$Pr(|P-\hat{P}|\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp{(-m\epsilon^{2}/2)}$ (16)
Proof:
According to the inequality for $L_{1}$ deviation of the empirical distribution [weissman2003inequalities], they conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\phi(\pi_{P})\epsilon^{2}/4)$, where $\phi(p)=\frac{1}{1-2p}\log\frac{1-p}{p}$ and $\pi_{P}=\max_{A\in\mathcal{A}}\min(P(A),1-P(A))$. The paper [weissman2003inequalities] point out that $0\leq\pi_{P}\leq 1/2,\forall P$. Based on this, we can derive the function $\phi(p)$ and find that $\phi(p)$ is monotonically decreasing on $(0,1/2)$. Therefore, we can get $\phi(p)\geq\phi(1/2)=2$. Then, we bring this result back and can conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\epsilon^{2}/2)$.
8.2 Missing Proofs
Theorem 8.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (17)
Proof:
$|J(\theta)-J^{*}(\theta)|$
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]-\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (18)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (19)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]$
$+\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (20)
$\leq\mathbb{E}_{m,x}[|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]|$
$+|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]|]$ (21)
$=\mathbb{E}_{m,x}[|V^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}|+|V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{m}|]$ (22)
$\leq\mathbb{E}_{m,x}[\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta)||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta))$
$+\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)||\pi(\cdot|s,Z^{*}(\cdot|x);\theta))]$ (23)
$\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (24)
Eq. (23) is the result of directly applying Lemma 8.1 and Eq. (24) is the result of directly applying Assumption 4.1. ∎
Theorem 8.4 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (25)
Proof:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ (26)
$=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (27)
$=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]]$ (28)
Denote $\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{i}(\cdot|x);\theta_{j})}}[R_{m}(s,a)]]$ as $G^{\theta_{j}}_{Z_{i}}$, then we can get the following derivation. Here, we only consider the value within the brackets.
$=\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (29)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (30)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (31)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (32)
$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))+\epsilon^{mutual}_{12}$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1}))$ (33)
$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}$
$-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (34)
Simplifying Eq. (34) we can get:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (35)
Theorem 8.5 (Monotonic performance improvement guarantee on training process).
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (36)
Here, $\xi\in[0,1]$ is a constant.
Proof:
Let $Z_{1}$ have already been trained by maximizing $I(Z;M)$ to some extent. Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ samples from the training dataset by maximizing $I(Z;M)$ to get the monotonic performance improvement guarantee.
We begin with the following inequality:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|+|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ (37)
where $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
According to Assumption 4.9, we have:
$|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (38)
Now, we move our focus on solving $|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $Z^{empirical}_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (39)
$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (40)
Then for a fixed $x$, with probability greater than $1-\xi$, we have:
$|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (41)
Hence, we can get:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (42)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ (43)
$k\geq\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (44)
Since we need the least number of samples that update the context encoder, we take $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$. ∎
Corollary 8.6 (Monotonic performance improvement condition for pre-training scheme).
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (45)
Proof: Denote $J^{pretrain}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the policy. Similarly, denote $J^{pretrain}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the policy.
$J^{pretrain}(\theta_{2})-J^{pretrain}(\theta_{1})$ (46)
$=J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (47)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (48)
$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))]+\epsilon^{mutual}_{12}$ (49)
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1}))]$ (50)
$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}]$ (51)
To get the monotonic performance improvement, if we consider the condition more strictly, we need the term within the expectation of Eq. (51) to be larger than 0. Hence, we can get:
$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{1}{2\kappa L_{z}}\epsilon^{mutual}_{12}=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (52)
8.3 Implementation Details
In this section, we report the implementation details of all the objectives in our main paper.
  • Contrastive-based [li2020focal]. We adopt the open-source code of FOCAL [^3] https://github.com/FOCAL-ICLR/FOCAL-ICLR/.
  • Reconstruction-based [li2024towards]. We obtain the source code of UNICORN from the authors and adopt UNICORN-0 as the reconstruction-based objective training framework.
  • Cross-entropy-based. Built upon the code base of FOCAL, cross-entropy-based only replaces the distance metric learning loss with cross-entropy loss.
To maintain fairness, we make sure all benchmark-irrelevant parameters are consistent, as shown in Table 1.
Table 1: Benchmark-irrelevant parameters setting in the training process.
training tasks 20
testing tasks 20
task training batch size 16
rl batch size 256
context size 1 trajectory
actor-network size [256, 256]
critic-network size [256, 256]
task encoder network size [64, 64]
learning rate 3e-4
Table 2: Benchmark-relevant parameters setting in the training process.
Configurations Reach Ant-Dir Button-Press Dial-Turn
dataset size 3e5 9e4 3e5 3e5
task representation dimension 5 5 5 5
For each environment, we make the benchmark-relevant parameters the same for each algorithm, as shown in Table 2.